# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 4.2 Unforgeability
In $\mathsf{Game}_{4}$, we construct an M-SIS adversary $\mathcal{B}$. Given M-SIS instance $\mathbf{A}$, $\mathcal{B}$ sets up the public key with $\mathbf{t}_{1}$ computed from $\mathbf{A}$ and runs $\mathcal{A}$ in $\mathsf{Game}_{4}$. If $\mathcal{A}$ outputs forgery $(\mu^{*},\sigma^{*})$, apply the forking lemma [PS00, BN06]: rewind and obtain two forgeries $(\mathbf{z}_{1},c_{1},\mathbf{h}_{1})$ and $(\mathbf{z}_{2},c_{2},\mathbf{h}_{2})$ with $c_{1}\neq c_{2}$ on the same $\mathbf{w}_{1}$. By verification:
$\mathsf{HighBits}(\mathbf{A}\mathbf{z}_{i}-c_{i}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}\text{ for }i\in\{1,2\}$
Define $\mathbf{v}=\mathbf{z}_{1}-\mathbf{z}_{2}$ and $\Delta c=c_{1}-c_{2}\neq 0$ (non-zero since $c_{1}\neq c_{2}$). Then $\mathbf{A}\mathbf{v}=\Delta c\cdot\mathbf{t}_{1}\cdot 2^{d}+\boldsymbol{\delta}$ where $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$. Rearranging:
$\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
This is an inhomogeneous M-SIS instance (SelfTargetMSIS [KLS18]): the residual $\boldsymbol{\delta}$ arises from rounding in $\mathsf{HighBits}$ and satisfies $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$, which is absorbed into the M-SIS norm bound. The extracted vector $(\mathbf{v},\Delta c)$ satisfies $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$. This is the standard reduction used in the Dilithium/FIPS 204 security proof. Thus: $\Pr[\mathsf{Win}_{4}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
@@ Conclusion.
Games 4 and the M-SIS reduction give $\Pr[\mathsf{Win}_{3.5}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$. Composing with the Rényi hop (Game $3\to 3.5$) and the birthday bound (Game $1\to 2$):
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
For ML-DSA-65 with $|S|\leq 17$: $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$, $n\ell=1280$), $\epsilon_{\mathsf{IH}}<10^{-30}$, and SHAKE-256 yields negligible $\epsilon_{\mathsf{PRF}}$, giving the stated bound.
@@ 4.3 Privacy
With Shamir nonce DKG, for any signing set $S$ with $|S|\geq T$ and any coalition $C\subset S$ with $|C|\leq T-1$: each honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy
$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
where $\mathsf{View}_{C}$ denotes the adversary’s view: corrupted shares, DKG evaluations, challenge $c$, the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$, and all signatures (current and prior). Note that $\mathsf{View}_{C}$ does not include individual $\mathbf{W}_{i}$ values, which are sent only to the coordinator (Remark 3.7). Since $\mathbf{w}_{1}$ is already implicit in any valid signature, it adds no information beyond $\sigma$ itself. For ML-DSA-65 ($n\ell=1280$, $\gamma_{1}=2^{19}$), the bound exceeds $1280\cdot(20-\log_{2}|S|)$ bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 1280\cdot\log_{2}9\approx 4058$ bits for $|S|\leq 17$. No computational assumptions are needed; the privacy is statistical.
Proof 4.5.
@@
Null space argument. The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$. The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
@@
Posterior constraint on $\mathbf{t}$. The coefficient $\mathbf{a}_{h,T-1}$ is sampled $R_{q}^{\ell}$-uniformly a priori. However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy the bounded-nonce constraint $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$. Since $\mathbf{n}(0)=\prod_{j\in C}(-x_{j})\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values. The posterior distribution of $\mathbf{t}$ is therefore not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of size $2\lfloor\gamma_{1}/|S|\rfloor+1$.
@@
Min-entropy bound. Since $\mathbf{n}(x_{h})=\prod_{j\in C}(x_{h}-x_{j})\neq 0\pmod{q}$ (because $q$ is prime and all evaluation points are distinct), multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$. The honest party’s evaluation $\mathbf{f}_{h}(x_{h})=\mathbf{g}(x_{h})+\mathbf{t}\cdot\mathbf{n}(x_{h})$ therefore inherits the same per-coefficient support size. The nonce share $\mathbf{y}_{h}=\sum_{j\in C}\mathbf{f}_{j}(x_{h})+\mathbf{f}_{h}(x_{h})$ shifts by a known constant, preserving support. The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again, yielding:
@@
$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
Remark 4.6 (Why SD =0=0 Is Unattainable).
@@
The bounded-nonce constraint ($\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$) restricts the free parameter $\mathbf{t}$ to $\approx 2\gamma_{1}/|S|$ values per coefficient, whereas $\mathbb{Z}_{q}$-uniformity would require $q\approx 2^{23}$ values. The per-coefficient statistical distance from uniform is at least $1-2\gamma_{1}/(|S|\cdot q)>0$ for all practical parameters. Achieving SD =0=0 would require either unbounded nonces (incompatible with rejection sampling) or adding pairwise masks on $\mathbf{z}_{i}$ with $|S|\geq T+1$. The min-entropy guarantee is nonetheless strong: for $|S|\leq 17$, nonce entropy exceeds secret key entropy by $\geq 5\times$, ensuring robust share privacy without computational assumptions.
@@
In Profiles P1 (TEE) and P3+ (trusted combiner), the protocol achieves response privacy with signing sets of size $|S|=T$ (minimum threshold). The two-honest assumption ($|S|\geq T+1$) is not needed: individual $\mathbf{W}_{i}$ values stay within the coordinator, and the min-entropy bound of Theorem 4.4 holds even with a single honest party. In Profile P2 (no coordinator), parties broadcast $\mathbf{W}_{i}$ directly; mask hiding (Lemma 4.9) protects these values under the condition $|S\setminus C|\geq 2$, which P2 already requires for r0-check share privacy.
@@
Over $K$ signing sessions with fresh nonce DKGs, the total min-entropy of honest party $h$’s responses $(\mathbf{z}_{h}^{(1)},\ldots,\mathbf{z}_{h}^{(K)})$ is at least $K\cdot n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$ bits (each session contributes independently). For $|S|\leq 17$, this exceeds $K\cdot 24{,}000$ bits, providing robust multi-session privacy. We recommend periodic key rotation per standard operational practice.
@@
Proof 4.10.
Since $|S\setminus C|\geq 2$, party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{nonce})$ for some honest $k^{\prime}\neq j$. This seed is unknown to $\mathcal{A}$, so by PRF security, the term is indistinguishable from uniform. Thus $\mathbf{V}_{j}=(\text{fixed})+(\text{uniform-looking})\stackrel{{\scriptstyle c}}{{\approx}}\text{uniform}$. The same argument applies to $\mathbf{W}_{j}$.
Note: Mask hiding requires $|S\setminus C|\geq 2$. In P1 and P3+, individual $\mathbf{W}_{i}$ are sent only to the coordinator (Remark 3.7), so mask hiding for $\mathbf{W}_{i}$ is only needed in P2 where parties broadcast directly. The response $\mathbf{z}_{i}$ achieves statistical privacy with even $|S\setminus C|=1$ via the nonce DKG (Theorem 4.4), independent of mask hiding.

### v5

@@ 4.2 Unforgeability
In $\mathsf{Game}_{4}$, we construct a SelfTargetMSIS adversary $\mathcal{B}$ [KLS18]. Given SelfTargetMSIS instance $(\mathbf{A},\mathbf{t}_{1})$ with $\mathbf{A}\in R_{q}^{k\times\ell}$ and $\mathbf{t}_{1}\in R_{q}^{k}$, $\mathcal{B}$ uses $(\mathbf{A},\mathbf{t}_{1})$ directly as the threshold scheme’s verification key components and runs $\mathcal{A}$ in $\mathsf{Game}_{4}$, answering ROM queries by programming $H$ uniformly at random, and recording all query-response pairs. $\mathcal{B}$ picks a uniformly random index $i^{*}\in[q_{H}]$ as the “target” ROM query. When $\mathcal{A}$ outputs a forgery $(\mu^{*},\mathbf{z}^{*},c^{*},\mathbf{h}^{*})$, $\mathcal{B}$ checks whether $c^{*}$ equals the response at query $i^{*}$; if not, $\mathcal{B}$ aborts (probability $1/q_{H}$ of success). When the guess is correct, verification gives:
$\mathsf{HighBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}^{*}$
The verification equation expands as:
$\mathbf{A}\mid-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{z}^{*}\\ c^{*}\end{bmatrix}=2\gamma_{2}\,\mathbf{w}_{1}^{*}+\mathbf{r}_{0}^{*}$
where $\mathbf{r}_{0}^{*}=\mathsf{LowBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$ satisfies $\|\mathbf{r}_{0}^{*}\|_{\infty}\leq\gamma_{2}$. The solution vector satisfies $\|(\mathbf{z}^{*},c^{*})\|_{\infty}\leq\max(\gamma_{1}-\beta,1)<\gamma_{1}$, solving the SelfTargetMSIS instance (known target $2\gamma_{2}\mathbf{w}_{1}^{*}$, LowBits residual $\mathbf{r}_{0}^{*}$) with the appropriate norm bound. This is the SelfTargetMSIS direct reduction used in the FIPS 204 security proof [KLS18]. Since $\Pr[\mathcal{B}\ \text{succeeds}]\geq\Pr[\mathsf{Win}_{4}]/q_{H}$, we have $\Pr[\mathsf{Win}_{4}]\leq q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
@@ Conclusion.
Games 4 and the M-SIS reduction give $\Pr[\mathsf{Win}_{3.5}]\leq q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$. Composing with the Rényi hop (Game $3\to 3.5$, joint divergence $(R_{2}^{\mathsf{vec,shift}})^{q_{s}}$) via Lemma Q.13 and the birthday bound (Game $1\to 2$):
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
For ML-DSA-65 with $|S|\leq 17$: $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{1280}\approx 1.0085$ (direct shift-invariance, Corollary Q.24); $\epsilon_{\mathsf{IH}}<10^{-30}$ (smooth Rényi tail parameter, Theorem 4.17); SHAKE-256 yields negligible $\epsilon_{\mathsf{PRF}}$. The security loss is $\approx 6.1\times 10^{-3}\cdot q_{s}$ bits, giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (non-vacuous for $q_{s}<16{,}000$; conservative bound in Remark 4.5; 96-bit figure assumes $q_{H}=1$, see Remark 4.6).
Remark 4.5 (Conservative Raccoon-Style Rényi Bound).
The analytical Raccoon-style formula (Lemma 4.2 of [Raccoon2024]) upper-bounds the per-coordinate Rényi divergence using $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+|S|^{4}\beta^{2}/(4\gamma_{1}^{2})\leq 1.003$ for $|S|\leq 17$, giving full-vector $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}\approx 2^{5.4}$ and a multi-session loss of $\approx 2.7\cdot q_{s}$ bits. This bound is vacuous for $q_{s}\gtrsim 36$ signing queries. Theorem 4.2 instead uses the direct FFT computation (Corollary Q.24, Theorem Q.25), which gives $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ per coordinate for $|S|=17$, yielding $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$—an improvement of ${\approx}440\times$. The conservative bound should not be used for concrete security estimates.
Remark 4.6 (The 96-Bit Baseline and $q_{H}$ Dependence).
The “96-bit baseline” in Theorem 4.2 arises from applying Cauchy-Schwarz to the Rényi transfer: $\sqrt{q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}}=2^{-96}$ when $\epsilon_{\mathsf{M\text{-}SIS}}=2^{-192}$ (NIST Level 3 hardness) and $q_{H}=1$. For $q_{H}$ random oracle queries the pre-Irwin-Hall security baseline is $\tfrac{1}{2}(192-\log_{2}q_{H})$ bits—e.g. $66$ bits for $q_{H}=2^{60}$. This $q_{H}$ dependence is a standard proof-technique artifact of ROM-based Fiat-Shamir proofs and is identical in the single-signer FIPS 204 security analysis [KLS18, FIPS204]; it is not specific to the threshold construction. The Irwin-Hall security loss of $0.006\cdot q_{s}$ bits (Corollary Q.24) is independent of $q_{H}$ and represents the additional cost introduced by the threshold nonce sharing above the single-signer baseline.
@@ 4.3 Privacy
With Shamir nonce DKG, for any signing set $S$ with $|S|\geq T$ and any coalition $C\subset S$ with $|C|\leq T-1$: each honest party’s nonce share $\mathbf{y}_{h}$ has conditional min-entropy
$H_{\infty}(\mathbf{y}_{h}\mid\mathsf{View}_{C}^{\mathsf{DKG}})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
where $\mathsf{View}_{C}^{\mathsf{DKG}}$ consists of: (1) the corrupted parties’ key shares; (2) nonce DKG evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ of the honest party’s polynomial at corrupted indices; and (3) cross-evaluations $\{\mathbf{f}_{j^{\prime}}(i)\}$ (for $j^{\prime}\in S\setminus C$, $j^{\prime}\neq h$, $i\in C$) of other honest parties’ polynomials. (Equivalently, as corrupted parties generate their own polynomials, the adversary’s view includes $\{\mathbf{f}_{j}\}_{j\in C}$ completely; the formal statement in Theorem Q.3 uses this equivalent conditioning.) The cross-evaluations $\{\mathbf{f}_{j^{\prime}}(i)\}$ from other honest parties to corrupted parties do not affect the min-entropy of $\mathbf{f}_{h}(x_{h})$: by Shamir’s security theorem, $|C|\leq T-1$ evaluations of any other degree-(T1)(T-1) polynomial $\mathbf{f}_{j^{\prime}}$ are independent of its (irrelevant) constant term, and none of these evaluations are of $\mathbf{f}_{h}$ itself. Additionally, since each $j\in C$ is corrupted, the adversary generated $\mathbf{f}_{j}$ and knows all evaluations $\mathbf{f}_{j}(x_{h})$ (sent privately to honest party $h$); these appear as a known additive constant $\sum_{j\in C}\mathbf{f}_{j}(x_{h})$ in the share $\mathbf{y}_{h}$, which preserves the support size (and hence min-entropy) of the free term $\mathbf{f}_{h}(x_{h})$.
For ML-DSA-65 ($n\ell=1280$, $\gamma_{1}=2^{19}$), the bound equals approximately $1280\cdot(20-\log_{2}|S|)$ bits (the exact formula $1280\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$ differs by <3<3 bits for $|S|\leq 17$)—over $5\times$ the short secret key entropy $H(\mathbf{s}_{1})\leq 1280\cdot\log_{2}9\approx 4058$ bits for $|S|\leq 17$. (Note: Shamir shares $\mathbf{s}_{1,h}=\mathbf{f}(h)$ for $h\neq 0$ are $\mathbb{Z}_{q}$-uniform and thus not short; the entropy comparison is against the underlying short secret $\mathbf{s}_{1}=\mathbf{f}(0)$.) No computational assumptions are needed; the bound is purely statistical.
Proof 4.8.
@@
Null space argument. The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space of dimension $T-|C|\geq 1$. In the worst case |C|=T1|C|=T-1, the null space is spanned by $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$ (the unique monic degree-(T1)(T-1) polynomial with roots at $\{x_{j}\}_{j\in C}$). For |C|<T1|C|<T-1, there are T|C|>1T-|C|>1 free parameters, which can only increase the adversary’s uncertainty; hence the stated bound is conservative and the proof below treats the hardest case |C|=T1|C|=T-1. The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
@@
Posterior constraint on $\mathbf{t}$. The highest-degree coefficient $\mathbf{a}_{h,T-1}$ is sampled $R_{q}^{\ell}$-uniformly a priori, so $\mathbf{t}=\mathbf{a}_{h,T-1}-\mathbf{g}_{T-1}$ is $\mathbb{Z}_{q}$-uniform per coefficient (the other high-degree coefficients are likewise $\mathbb{Z}_{q}$-uniform but are already determined by the adversary’s evaluations in the |C|=T1|C|=T-1 case). The constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy the bounded-nonce constraint $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$. Since $\mathbf{n}(0)=\prod_{j\in C}(-x_{j})\neq 0\pmod{q}$, multiplication by $\mathbf{n}(0)$ is a bijection on $\mathbb{Z}_{q}$; the constraint selects exactly $2\lfloor\gamma_{1}/|S|\rfloor+1$ values per coefficient of $\mathbf{t}$. Because the prior is $\mathbb{Z}_{q}$-uniform, conditioning on membership in this subset yields a uniform posterior over the $2\lfloor\gamma_{1}/|S|\rfloor+1$ valid values per coefficient.
@@
Min-entropy bound. Since $\mathbf{n}(x_{h})=\prod_{j\in C}(x_{h}-x_{j})\neq 0\pmod{q}$ (because $q$ is prime and all evaluation points lie in $\{1,\ldots,N\}$ with $N\leq q-1$, so they are distinct nonzero elements of $\mathbb{Z}_{q}$), multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$. The honest party’s evaluation $\mathbf{f}_{h}(x_{h})=\mathbf{g}(x_{h})+\mathbf{t}\cdot\mathbf{n}(x_{h})$ therefore inherits the same per-coefficient support size. The full nonce share is
$\mathbf{y}_{h}\;=\;\underbrace{\sum_{j\in C}\mathbf{f}_{j}(x_{h})}_{\text{known to }\mathcal{A}}\;+\;\mathbf{f}_{h}(x_{h})\;+\;\underbrace{\sum_{j\in S\setminus C,\,j\neq h}\mathbf{f}_{j}(x_{h})}_{\text{unknown to }\mathcal{A};\;\text{adds further entropy}}.$
The first sum is a known constant (from corrupted parties’ polynomials). The third sum is unknown (each $\mathbf{f}_{j}(x_{h})$ with $j\neq h$ honest was sent only to party $h$), and the individual terms $\mathbf{f}_{j}(x_{h})$ for distinct honest parties $j,j^{\prime}\in S\setminus(C\cup\{h\})$ are independent conditional on the adversary’s view: each polynomial $\mathbf{f}_{j}$ is sampled independently, and the adversary observes only evaluations $\{\mathbf{f}_{j}(i)\}_{i\in C}$ with $i\neq x_{h}$ (since $h\notin C$); these $|C|\leq T-1$ evaluations determine a (T|C|)(T-|C|)-dimensional affine subspace of polynomials, within which the evaluation $\mathbf{f}_{j}(x_{h})$ remains uniformly distributed over a set of size $q^{n\ell(T-|C|)}$ (the null space dimension). Since the adversary’s observations of different honest polynomials involve disjoint evaluation sets (each $\mathbf{f}_{j}$ evaluated only at corrupted parties’ points $\{i\in C\}$, which do not include $x_{h}$), the conditional distributions of $\mathbf{f}_{j}(x_{h})$ and $\mathbf{f}_{j^{\prime}}(x_{h})$ are independent, so the third sum can only increase entropy. Using only the entropy contribution from $\mathbf{f}_{h}(x_{h})$ as a conservative lower bound (discarding the additional entropy from the third sum):
@@
$H_{\infty}(\mathbf{y}_{h}\mid\mathsf{View}_{C}^{\mathsf{DKG}})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
Remark 4.9 (Key Privacy of the Aggregated Signature).
Theorem 4.7 establishes privacy of the nonce shares $\mathbf{y}_{h}$—a purely statistical guarantee requiring no computational assumptions. Key privacy of the aggregated signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$—whether $\sigma$ reveals $\mathbf{s}_{1}$ beyond the public key $\mathbf{t}_{1}$—is a separate, stronger property not established in this work. It remains an open problem in the broader ML-DSA literature, unresolved even for single-signer FIPS 204: the equation $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ with sparse challenge $c$ does not directly reduce to a standard Module-LWE instance, and no formal reduction is known. All threshold variants inherit this open problem from the single-signer setting.
Remark 4.10 (Why SD =0=0 Is Unattainable (with Derivation)).
The bounded-nonce constraint ($\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$) restricts the free parameter $\mathbf{t}$ to $\approx 2\gamma_{1}/|S|$ values per coefficient, whereas $\mathbb{Z}_{q}$-uniformity would require $q\approx 2^{23}$ values.
Derivation of SD lower bound. Let $P_{\mathsf{post}}$ be the posterior distribution of a single coefficient of $\mathbf{f}_{h}(x_{h})$ given the adversary’s view (Theorem 4.7 shows this is uniform over $M=2\lfloor\gamma_{1}/|S|\rfloor+1$ values), and let $P_{\mathsf{unif}}$ be the uniform distribution over $\mathbb{Z}_{q}$. The statistical distance is:
$\mathsf{SD}(P_{\mathsf{post}},P_{\mathsf{unif}})$ $=\tfrac{1}{2}\sum_{x\in\mathbb{Z}_{q}}\bigl|P_{\mathsf{post}}(x)-P_{\mathsf{unif}}(x)\bigr|$
$=\tfrac{1}{2}\Bigl(M\cdot\bigl|\tfrac{1}{M}-\tfrac{1}{q}\bigr|+(q-M)\cdot\tfrac{1}{q}\Bigr)$
$=\tfrac{1}{2}\Bigl(\bigl|1-\tfrac{M}{q}\bigr|+\tfrac{q-M}{q}\Bigr)$
$=1-\tfrac{M}{q}\;=\;1-\tfrac{2\lfloor\gamma_{1}/|S|\rfloor+1}{q}>0$
For ML-DSA-65 with $\gamma_{1}=2^{19}$, q=8,380,417q=8{,}380{,}417, and $|S|=17$: $M=2\lfloor 2^{19}/17\rfloor+1=61{,}681$, so $\mathsf{SD}=1-M/q\approx 1-2{\cdot}2^{19}/(17{\cdot}q)\approx 1-2^{20}/(17{\cdot}2^{23})=1-2^{-3}/17\approx 0.9926>0$ (very far from zero).
@@
Achieving SD =0=0 would require either unbounded nonces (incompatible with rejection sampling) or adding pairwise masks on $\mathbf{z}_{i}$ with $|S|\geq T+1$ (Two-Honest assumption). The min-entropy guarantee is nonetheless strong: for $|S|\leq 17$, nonce entropy exceeds secret key entropy by $\geq 5\times$, ensuring robust share privacy without computational assumptions.
@@
In Profiles P1 and P3+, the protocol achieves nonce share privacy with $|S|=T$. In Profile P2, mask hiding (Lemma 4.15) applies under $|S\setminus C|\geq 2$.
Proof 4.12.
[…]
@@
Let $M_{S}=2\lfloor\gamma_{1}/|S|\rfloor+1$. Over $K$ sessions with fresh nonce DKGs, total nonce min-entropy is at least $Kn\ell\cdot\log_{2}M_{S}$ bits. For $|S|\leq 17$, this exceeds $K\cdot 20{,}000$ bits. Key privacy across sessions is computational (multi-sample bounded-distance decoding under Module-LWE).
Proof 4.14.
Each signing session uses a fresh nonce DKG with independent randomness, so the free parameters $\mathbf{t}^{(1)},\ldots,\mathbf{t}^{(K)}$ are independent. The min-entropy of a product of independent distributions equals the sum: $H_{\infty}(\mathbf{y}_{h}^{(1)},\ldots,\mathbf{y}_{h}^{(K)}\mid\mathsf{View}_{C})=\sum_{k}H_{\infty}(\mathbf{y}_{h}^{(k)}\mid\mathsf{View}_{C}^{(k)})$. The adversary accumulates $K$ equations $\mathbf{z}_{h}^{(k)}=\mathbf{y}_{h}^{(k)}+c^{(k)}\cdot\mathbf{s}_{1,h}$ sharing the same $\mathbf{s}_{1,h}$; after sufficiently many sessions, $\mathbf{s}_{1,h}$ is uniquely determined by the overdetermined linear system $\mathbf{z}_{h}^{(k)}=\mathbf{y}_{h}^{(k)}+c^{(k)}\cdot\mathbf{s}_{1,h}$ over $\mathbb{Z}_{q}$ (as in single-signer ML-DSA). Key privacy across multiple sessions is therefore computational: extracting $\mathbf{s}_{1,h}$ from these equations amounts to a multi-sample bounded-distance decoding problem, which is hard under the Module-LWE assumption.
@@
Proof 4.16.
Since $|S\setminus C|\geq 2$, party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{session\_id}\|\mathsf{dom})$ for some honest $k^{\prime}\neq j$, where $\mathsf{session\_id}$ is the unique per-signing-session identifier (e.g., derived from the preprocessed nonce commitment $\mathsf{Com}_{j}$) and $\mathsf{dom}$ is the fixed domain label for $\mathbf{V}$ or $\mathbf{W}$. Because $\mathsf{seed}_{j,k^{\prime}}$ is unknown to $\mathcal{A}$ and $\mathsf{session\_id}$ is fresh each session, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{session\_id}\|\mathsf{dom})$ is computationally indistinguishable from a fresh uniform element of $R_{q}^{k}$ by PRF security. Thus $\mathbf{V}_{j}=(\text{fixed})+(\text{uniform-looking})\stackrel{{\scriptstyle c}}{{\approx}}\text{uniform}$. The same argument applies to $\mathbf{W}_{j}$.
Note: Mask hiding requires $|S\setminus C|\geq 2$. In P1 and P3+, individual $\mathbf{W}_{i}$ are sent only to the coordinator (Remark 3.9), so mask hiding for $\mathbf{W}_{i}$ is only needed in P2 where parties broadcast directly. The nonce share $\mathbf{y}_{i}$ achieves statistical privacy with even $|S\setminus C|=1$ via the nonce DKG (Theorem 4.7); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; see Remark 4.9), independent of mask hiding.